OpenStudy (nubeer):

find all roots of z^5+2z^3+6z^2+12 =0 using demoivre theroem anyone plz?

OpenStudy (lgbasallote):

the demoivre theorem i am familiar with is the one in Calculus. Is that the one you're pertaining to?

OpenStudy (nubeer):

yes its in caculus.

OpenStudy (mathmate):

You'll see why when you factor the left-hand side of the equation to (z^2+2)*(z^3+6) = 0

OpenStudy (nubeer):

ahh yes i think this is the factor this helps alot.. can u tell me how u got that... dont tell me its again hit and trial

OpenStudy (mr.math):

\[(z^5+2z^3)+(6z^2+12)=z^3(z^2+2)+6(z^2+2)=(z^3+6)(z^2+2).\]

OpenStudy (nubeer):

well thanks .. this way is also a good one.. can u tell me any website whichhave questions like these with solutions.

OpenStudy (mr.math):

This was just a simple factorization, You can search factorization techniques. You still need to solve \((z^3+6)(z^2+2)=0\), which implies \(z^2+2=0\) or \(z^3+6=0\).

OpenStudy (mr.math):

\(z^2+2=0\) is very easy to solve and you it gives the roots \(z=\pm \sqrt{2}i\).

OpenStudy (mr.math):

You can use the demoivre theorem to solve \(z^3+6\), (i.e find the third roots of \(-6\)).

OpenStudy (nubeer):

so this last term will give us three roots?

OpenStudy (mr.math):

Yep, The notes vishal kothari gave are very good. You should probably read them.

OpenStudy (nubeer):

ya i read them all.. they are helpful.. but still was kinda confused of how getting these roots.. but thank you.. do u know any website for more questions about this? with solutions?

OpenStudy (mr.math):

I don't, but you can search the internet for solutions of polynomials over complex numbers. I would look for you, but I have to leave now.

OpenStudy (anonymous):

put z=e^ix =cosx+isinx then proceed

OpenStudy (nubeer):

@Mr.Math . thank you very much have a nice day :)

OpenStudy (mr.math):

You're welcome!